OpenStudy (anonymous):

the dimensions of a rectangle are such that its length is 3 in more than its width. If the lenght decreased by 1 in. the area would be increased by 236in^2. What are the lenght and width of the rectangle?

OpenStudy (dumbcow):

A = x(x+3) wait is that a typo? if length is small the rectangular should have a smaller area?

OpenStudy (anonymous):

Typo: the dimensions of a rectangle are such that its length is 3 in more than its width. If the length doubled and if the width decreased by 1 in. the area would be increased by 236 in^2 . What are the length and width of the rectangle?

OpenStudy (dumbcow):

ahh that makes more sense A=x(x+3) --> A = x^2 +3x and A+236 = (x-1)*2(x+3) --> A+236 = 2x^2+4x-6 solve by substitution

OpenStudy (anonymous):

still not understand honestly

OpenStudy (dumbcow):

x is width, x+3 is length , A is area double length is 2*(x+3) , width -1 is just x-1 replace "A" with "x^2 +3x" into 2nd equation set equal to 0, then solve for x using quadratic formula

OpenStudy (anonymous):

Thanks =D

OpenStudy (dumbcow):

no problem